Cullumber, Inc. acquired 20% of Marigold Corporation's voting stock on January 1, 2021 for $930000. During 2021, Marigold earned $400000 and paid dividends of $248000. Cullumber's 20% interest in Marigold gives Cullumber the ability to exercise significant influence over Marigold's operating and financial policies. During 2022, Marigold earned $524000 and paid cash dividends of $174000 on April 1 and $174000 on October 1. On July 1, 2022, Cullumber sold half of its stock in Marigold for $684000 cash. What should the gain be on sale of this investment in Cullumber's 2022 income statement?
Answer #1

Calculate value of investment

Initial investment 930000
Add: Net income (2021) (400000*20%) 80000
Less; Dividend (248000*20%) -49600
Ending investment 2021 960400
Add: Net income (524000*20%) 104800
Less: Dividend (174000*20%) -34800
Value of sale 1030400

Value of Half stock = 1030400/2 = 515200

Sale price = 684000

Gain on Sale of stock = 684000-515200 = 168800